Km and Vmax homework problem Given: [S] M n mmoles/liter-min 13.8 16.0 8.35 (E-6) 1.00 (E-5) 1.25 (E-5) 1.67 (E-5) 2.0 (E-5) 2.5 (E-5) 3.3 (E-5) 5.0 (E-5) 1.0 (E-4) 2.0 (E-4) 19.1 23.8 26.7 30.8 36.2 44.5 57.2 66.7 Find: Km and Vmax using at least three methods.
Answer #1

[S] M V M moles/liter-min

1 16

1.25 19.1

1.67 23.8

2 26.7

2.5 30.8

3.3 36.2

5 44.5

1/[s] M-1 1/V Mmol-1 liter-min

1 0.0625

0.8 0.052356

0.598802 0.042017

0.5 0.037453

0.4 0.032468

0.30303 0.027624

0.07 0.049x+ 0.012 0.06 0.05 0.04 0.03 0.02 0.01 Series1 _Linear (Series1) 0.5 1.5

We know that, (Linweaver and bulk plot)

V= (Vmax [s])/(Km+[s])

1/v =( Km/Vmax) 1/([s]) + 1/Vmax

From graph y=0.049x+.0012

Hence, 1/Vmax=.0012Mmol-1 liter-min

Vmax= 83.33 M moles/liter-min

Now,Km/Vmax = 0.049

Km= (0.049*83.33) = 4.083M

v